OpenStudy (anonymous):

4x^2 =8 solve using radicals,rstionalize all denominators express complex numbers in terms of i or use a comma to separate answers if needed

OpenStudy (anonymous):

then what would the x intercepts be

OpenStudy (anonymous):

y=4x^2 so by this fomrula y=mx+b we will have the slope=4 and y-intercept is =0

OpenStudy (anonymous):

what would be the first part of the question?

OpenStudy (anonymous):

4x^2 -8 = 0 is a quadratic (parabola) with roots plus or minus sqrt 2.